OpenStudy (anonymous):

Steve traveled 200 miles at a certain speed. Had he gone 10mph faster, the trip would have taken 1 hour less. Find the speed of his vehicle

OpenStudy (anonymous):

Distance = Rate x Time Distance is equal in each case 200 miles. D = RT = (R+10) x (T-1) Answer is in case 1 the car is driving 5 hours at 40 mph In case 2 -- the car is going 10 mph greater which is 50 mph, and gets there in 4 hours 5hrs x 40 mph = 200 miles 4 hrs x 50 mph = 200 miles

OpenStudy (callisto):

similar to the above. as the distance travelled is thte same, (s+10)(t-1) = st 10t-s-10 =0 10t-s=10 ---(1) since s=d/t, t=200/s --(2) Sub(2) into (1) 10(200/s) -s =10 s^2 +10s - 2000 =0 s=40mph or s=-50mph (rejected) therefore, speed required is 40 mph

OpenStudy (anonymous):

thank you